Please, can any one help me solve this difficult problem?? I would deeply appreciate it. Thank you!!
4(x+2)^2+320=0

Answer by edjones(8007)

4(x+2)^2+320=0
4(x^2+4x+4)+320=0
4x^2+16x+16+320=0
4x^2+16x+336
4(x^2+4x+84)=0
x^2+4x =-84
x^2+4x+4=-84+4 complete the square
(x+2)^2=-80
x+2=+-sqrt(-80)
x=-2+-sqrt(16*-5)
x=-2+-4i*sqrt(5)
Ed